Adjusted 3D Family Fireworks Trial Balance January 31, 2021 Account Title Credit $ Debit 89,400 14,600 1,630 Cash Accounts ReOn January 1, 2021, the general ledger of 3D Family Fireworks includes the following account balances: Credit $ 25, 900 14, 6Journal entry worksheet 10 Record the closing entry for revenue. Note: Enter debits before credits. Credit Date General JournJournal entry worksheet < ..... 6 0 8 9 10 11 12 Record the closing entry for expenses. Note: Enter debits before credits. Gewill then populate the balances in those accounts from the adjusted trial balance. Post-closing - 3D Family Fireworks Income3D Family Fireworks Balance Sheet January 31, 2021 Assets Liabilities Current Assets Current Liabilities: 01 0 Total CurrentEnter your Accounts Receivable turnover value in 1 decimal place and Ratio of Allowance for Uncollectible Accounts in Whole A
Answer #1
Account Titles Debit Credit
12 Service Revenue $       1,19,500
Interest Revenue $                125
Retained Earnings $         1,19,625
(To close revenue accounts)
13 Retained Earnings $          85,580
Supplies expense $              2,750
Salaries Expense $            66,900
Utilities expense $            14,700
Bad Debt Expense $              1,230
(To close expense accounts)
Income Statement
Revenues
Service Revenue $       1,19,500
Interest Revenue $                125
Total Revenues $         1,19,625
Expenses
Supplies expense $             2,750
Salaries Expense $          66,900
Utilities expense $          14,700
Bad Debt Expense $             1,230
Total Expenses $            85,580
Net Income $            34,045
Statement of Retained Earnings
Beginning Balance $          33,400
Add : Net income $          34,045
Less Dividend $                   -  
Ending Balance $          67,445
Balance Sheet
Assets
Current Assets
Cash $          89,400
Accounts Receivable $          14,600
Less Allowance for Uncollectible accounts $           -1,630
Notes Receivable $          30,000
Interest Receivable $                125
Supplies $                750
Total Current Assets $        1,33,245
Non Current Assets
Land $            78,000
Total Assets $        2,11,245
Liabilities & Stockholder's Equity
Current Liabilities
Accounts Payable $             3,300
Salaries Payable $          34,500
Total Current Liabilities $            37,800
Stockholder's Equity
Common Stock $       1,06,000
Retained Earnings $          67,445
Total Stockholder's Equity $        1,73,445
Total Liabilities & Stockholder's Equity $        2,11,245

Accounts Receivable Turnover = Net Credit Sales / Accounts Receivable
= 82400/14600 = 5.64

True, the company is collecting more efficiently

Ratio of Allowance for Uncollectible accounts to accounts receivable = $1630/14600 = 11.16%

Beginning of year ratio = $2800/14600 = 19.18%

Improving conditions, since ratio declines